OpenStudy (anonymous):

If f(x)= (x+3)/3, then what is f^-1(x)???

OpenStudy (anonymous):

sorry, whats f^ is it a differential sign ?

OpenStudy (ajprincess):

Let us assume f^-1(x)=y Then f(y)=x Also f(y)=y+3/3 y+3/3=x y+3=3x y=3x-3 Hence f^-1(x)=3x-3 Is it clear?

OpenStudy (anonymous):

3x-3

OpenStudy (anonymous):

thank you aj & snap :)

OpenStudy (ajprincess):

U r welcome.:)